Suppose b,c belongs to R. Define map T : R3 --> R2 by
T.(x,y,z)= (2x - 4y + 3z +b, 6xcxyz)
Show that T is linear if and only if b = c = 0.
1
Expert's answer
2020-06-15T13:11:57-0400

Let us consider "T(x_1+x_2,y_1+y_2,z_1+z_2)." It should be equal to "T(x_1,y_1,z_1) + T(x_2,y_2,z_2)" .

"T(x_1+x_2,y_1+y_2,z_1+z_2) = \big(2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+b, 6c(x_1+x_2)^2(y_1+y_2)(z_1+z_2)\big)."

"T(x_1,y_1,z_1) + T(x_2,y_2,z_2) = (2x_1-4y_1+3z_1+b,6cx_1^2y_1z_1) + (2x_2-4y_2+3z_2+b,6cx_2^2y_2z_2 ) = \big(2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+2b, 6c(x_1^2y_1z_1+x_2^2y_2z_2)\big)."


Therefore, "2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+b = 2(x_1+x_2)-4(y_1+y_2) + 3(z_1+z_2)+2b" for every "x_1,x_2,y_1,y_2,z_1,z_2\in\R." It is true if and only if "2b=b \; \Rightarrow b = 0."


Also it should be true that "6c(x_1+x_2)^2(y_1+y_2)(z_1+z_2) = 6c(x_1^2y_1z_1+x_2^2y_2z_2)" for every "x_1,x_2,y_1,y_2,z_1,z_2\in\R." Let, for example, "x_1=y_1=z_1 = 1, \; x_2=y_1=z_1=2" , therefore

"6c\cdot81= 6c\cdot65 \Rightarrow c = 0." So "c" can be only equal to 0.


So we conclude that if T is linear, b=0 and c = 0. Now we should prove that for "b=c=0" T is linear.


"T(x,y,z) = (2x - 4y + 3z, 0)" . For every "x_1,x_2,y_1,y_2,z_1,z_2\in\R \; \;" "T(x_1+x_2,y_1+y_2,z_1+z_2) = (2(x_1+x_2)-4(y_1+y_2)+3(z_1+z_2),0) = (2x_1-4y_1+3z_1,0) + (2x_2-4y_2+3z_2,0) = T(x_1,y_1,z_1) + T(x_2,y_2,z_2),"

so T is linear.